### v1

# Depends-Kotlin: A Cross-Language Kotlin Dependency Extractor

Abstract.
Since Google introduced Kotlin as an official programming language for developing Android apps in 2017, Kotlin has gained widespread adoption in Android development. However, compared to Java, there is limited support for Kotlin code dependency analysis, which is the foundation to software analysis. To bridge this gap, we developed Depends-Kotlin to extract entities and their dependencies in Kotlin source code. Not only does Depends-Kotlin support extracting entities’ dependencies in Kotlin code, but it can also extract dependency relations between Kotlin and Java. The extraction of such cross-language dependencies can help developers understand the migration process from Java to Kotlin. Additionally, we used a Java project with confirmed dependencies as a benchmark and converted this project to two projects: Kotlin-only and a combination of Kotlin and Java. The dependencies in these two projects were then extracted using our tool. The consistency observed among dependency relations in all three projects confirms the accuracy of Depends-Kotlin. Furthermore, the performance of Depends-Kotlin was assessed using another three projects of varying sizes. The source code of Depends-Kotlin and the dataset used in this demo paper have been uploaded to https://github.com/XYZboom/depends-kotlin. We also provided a screencast presenting Depends-Kotlin https://youtu.be/daZuXOwn1Ls.
Kotlin, Code Dependency, Cross-Language
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/xxxxxxx.xxxxxxxconference: FSE’24: the 32nd ACM Symposium on the Foundations of Software Engineering (FSE ’24) 2024; November 15–19, 2024; Porto de Galinhas, Brazilbooktitle: Proceedings of the 32nd ACM Symposium on the Foundations of Software Engineering (FSE ’24), November 15–19, 2024, Porto de Galinhas, Brazilisbn: 978-1-4503-9999-9/18/06
1. Introduction
The dependency relations among entities in the source code form the foundations of software architecture analysis, including architecture recovery, architecture anti-pattern detection, architecture quality evaluation, and more (Mo et al., 2019; Cui et al., 2019; Xiao et al., 2022). Since Google introduced Kotlin as an official programming language for developing Android apps in 2017, Kotlin has gained widespread adoption in Android development. According to recent empirical studies, a significant of Android apps have been continuously migrated from Java to Kotlin (Ardito et al., 2020; Mateus and Martinez, 2020). However, compared to dependency analysis tools that support Java, such tool support for Kotlin is limited. To our knowledge, many well-known dependency analysis tools that support the Java language, such as Structure 101 (str, [n. d.]), Understand (und, [n. d.]), and DV8 (dv8, [n. d.]), currently do not offer support for the Kotlin language.
Kotlin dependency resolution faces two challenges. First, Kotlin was introduced by Google as “concise, expressive, and designed to be type and null-safe”. It contains a lot of syntactic sugar to ensure these features, thereby increasing the difficulty of resolving entity types and dependencies. Consider the following Kotlin code: Line 1 declares a class named Bar with a property x of type Int (property in Kotlin is similar to field in Java). Line 4 declares a high-level function named calculate that takes a lambda expression as its parameter. As shown in this line, the lambda takes a Bar type (called receiver type in Kotlin), and returns an Int. Line 7 declares another class named Foo, also with a property x of type Int. Line 8 declares calculateInFoo, which is a member function of the Foo class. It calls the calculate function in Line 9. Since the lambda parameter takes type Bar, the add function invoked in the lambda accesses the x property of the current Bar instance, not the Foo instance. The use of lambdas with receiver types as input allows concise syntax but it increases the difficulty of dependency resolution. In order to resolve this dependency (calculateInFoo use Bar.x), we need to locate calculate’s parameters and further trace them down to the correct type and properties. This is different from Java dependency analysis. Furthermore, Kotlin has its own dependency types, such as delegate and extension, which are supported by our tool and will be discussed in more detail in Section 2.
1class Bar(val x: Int)
2// Declaration of class Bar with the property x
3
4fun calculate(param: Bar.() -> Int) {}
5/* Function ’calculate’ takes a lambda with a receiver type Bar as its parameter.*/
6
7class Foo(val x: Int) {
8    fun calculateInFoo() {
9        calculate { add(x) }
10        // x in add(x) here is actually Bar.x
11    }
12}
Second, Kotlin is designed to be fully interoperable with Java and can run on the JVM, which makes it easy for developers to continuously migrate Java code to Kotlin. The following example includes a Kotlin class BarKotlin and a Java class FooJava. Lines 4-5 demonstrate that the Java class FooJava accepts and interacts with the Kotlin class BarKotlin by invoking a getX() method of the Kotlin class. However, in the Kotlin code, there is no explicit getX() method, as the getter method is implicitly generated by JVM for the class property. If we analyze only static code, dependencies between these two languages cannot be recognized. Considering that a significant percentage of Android apps in migration involve both Java and Kotlin code (Ardito et al., 2020; Mateus and Martinez, 2020), a tool for resolving Kotlin dependencies should not only address dependency relations in Kotlin code but also handle dependencies between Java and Kotlin code.
1class BarKotlin(val x: Int)
2public class FooJava {
3    public static void func(BarKotlin bar) {
4        System.out.println(bar.getX());
5    }
6}
To tackle these two challenges, we propose Depends-Kotlin, a cross-language Kotlin dependency extractor. Depends-Kotlin is based on the Depends framework, an open-source project designed for code dependency analysis (dep, [n. d.]; Jin et al., 2019). We enhanced Depends to support the Kotlin language by performing multi-round inference for getting Kotlin’s specific types and dependency relations. Additionally, we addressed cross-dependency relations between Kotlin and Java by refactoring the architecture of the original Depends framework and filling in the implicit code of Kotlin properties, which is handled by JVM and not shown in the source code.
To validate Depends-Kotlin’s accuracy, we manually converted a sample project written in Java into two separate projects: one exclusively containing only Kotlin code and the other featuring a combination of Kotlin and Java code. Using confirmed dependencies of Java code as the ground truth, we show that Depends-Kotlin can capture entities and their dependency relations accurately in both Kotlin code and Kotlin-Java scenarios. Additionally, we assessed Depends-Kotlin’s performance by analyzing the dependencies of another three Java-Kotlin projects with varing sizes.
2. Methodology
2.1. Framework
Figure 1 shows the architecture of Depends-Kotlin. We adopted a plug-in architecture (Wolfinger et al., 2006) and decomposed the original Depends architecture into two parts: Depends-Core (the main framework) and Depends-Lang (plug-ins). Depends-Lang manages the entity parsing for specific languages, while Depends-Core handles the resolution of dependencies for the parsed entities. This architecture refactoring decision aims to enable the handling of cross-language dependencies, which is not supported by the original Depends architecture. To this end, we added in Depends-Core a Language Register module, which leverages Java Service Provider Interface (SPI) mechanism and allows registering different programming languages simultaneously. For a specific language, its language processor needs to extend the interface and has been registered in the SPI file. We also modified the existing Relation Resolver module in Depends-Core to generate the implicit code, specifically the getter/setter code for Kotlin properties mentioned in Section  1. This modification helps the detection of Java-Kotlin interactions.
In order to process Kotlin’s unique syntax and resolve its dependency types, we modified the existing Expression Analyser and Relation Resolver modules in Depends-Core. Specifically, We analyzed Kotlin’s extension functions and properties in Expression Analyser. We also identified the scope of Kotlin’s functions by adding related code as the context in Expression Analyser. Furthermore, we adapted Relation Resolver to handle dependencies with built-in types and new unique dependency types in Kotlin. For example, as method members of built-in types in Java always return built-in types, the original Depends chose to focus on the dependencies of analyzed source files and ignored built-in types. However, due to Kotlin extension functions, built-in types can return any types. This can greatly impact dependencies in the analyzed source code. Consequently we modified the Relation Resolver module to resolve such extension type.
For Kotlin’s entity parsing, we adopted the same logic of Java-parser in the original Depends framework. We used Antlr v4 to generate an Abstract Syntax Tree (AST) parser from a Kotlin grammar, which is provided on the official Kotlin website. This module is named as Kotlin Processor as shown in Figure  1. Subsequently, the Kotlin EntityTree Builder module store parsed AST nodes into concrete entities, such as files, packages, types, expressions, functions, and properties. In this step, some expressions’ types can be resolved in advance, which can reduce subsequent workload. As previously mentioned, the parser module of a specific language can work as a plug-in to the Depends-Core framework, enabling the analysis of other languages in the future.
Figure 1. Plug-in architecture of Depends-Kotlin.
2.2. Entity and Relation Resolution
To achieve accuracy with a good performance, we followed the same logic of the original Depends framework and performed multi-round inference for getting entities and dependency relations. An entity is represented by a tuple (id, name, entityType, context) and a relation is represented by (sourceId, targetId, dependencyType, weight). For simple dependency relations, such as extend, function parameters, function return and delegate, such entities and dependency relations are stored temporarily when analyzing expressions. For dependency relations, such as member access and member function call on expressions of unknown types, we conducted type inference. The basic idea is to deduct types from what is known to what is unknown. Any operation on an expression with a known type, such as a member function call, will yield the next expression with a known type. This process is recursively repeated until the entire expression’s type inference is completed. At this point, all entities and their dependencies should have been resolved.
It is worth mentioning that an additional search is required to resolve Kotlin extension functions during this process, as the scope of the expression containing extension functions needs to be identified. While traversing the syntax tree, if a function has a receiver type, it is marked as an extension function and the extended type is the receiver type. After the type system processing is complete, extension functions can be located based on their marks during the traversal of the entity tree. The extended types and the extension relationship of the function can then be recorded.
Our tool framework also automatically generates implicit code for Kotlin, and it distinguishes Java and Kotlin entities by assigning them different labels. When resolving expressions involving Kotlin-Java interactions, such cross-language relations can be easily detected. Table 1 lists the dependency types currently supported by Depends-Kotlin. The first six dependency relations can be observed in both Java-only and Kotlin-only projects. The delegate and extension relations are exclusive to Kotlin projects. The last two, JDependsOnK and KDependsOnJ, can be observed in projects containing dependencies between Java and Kotlin code.
Table 1. Relation types supported by Depends-Kotlin.
Relation Description
Import a file imports another class, enum, static method
Extend a class extends a parent class or an interface
Call a statement in a method invokes another method
Use a method access a variable in its scope
Parameter a method use another type as its parameter
Return a method returns another type
Delegate a class delegates another class
Extension a method is an extension of a class
JDependsOnK Java code depends on Kotlin entities
KDependsOnJ Kotlin code depends on Java entities
3. Evaluation
As Depends-Kotlin is designed for Kotlin and Kotlin-Java dependency resolution, we conducted an evaluation to assess its capability in resolving dependency relations in both Kotlin and Kotlin-Java scenarios. The evaluation focused on two perspectives: accuracy and performance.
3.1. Accuracy Verification
We selected a Java project with confirmed dependencies as the ground truth. Subsequently, we converted the Java project to Kotlin-only and Kotlin-Java projects. We utilized Depends-Kotlin to extract dependencies in these (partially) converted projects. As these three projects provide the same features, they are supposed to have similar entities and dependencies. If the extracted dependencies in the converted projects align with the ground truth, it proves that Depends-Kotlin can accurately capture Kotlin and Kotlin-Java dependency relations. If not, we should be able to explain the discrepancies, which may be due to the syntax difference between Java and Kotlin. The Java project we chose is implementation of a maze game (Cai et al., 2013) and in this game a maze comprises multiple rooms, each containing walls and a door leading to another room. The initial source code, named as ‘maze-java’, is written in Java and consists of 21 classes. The dependencies of ‘maze-java’ are extracted using the original Depends tool and have been confirmed as the ground truth.
We first converted all Java source code in ‘maze-java’ to Kotlin source code, named ‘maze-kotlin’. This process was automatically conducted by the conversion tool in IntelliJ IDEA. Next, we converted the source code of certain Java classes (e.g., data, utility) from ‘maze-java’ to Kotlin, resulting in ‘maze-kotlin-java’. The rationale behind choosing data/utility classes is to simulate real-world Java to Kotlin migration processes within a project  (Mateus et al., 2023). In this process, we converted 4 classes: Rectangle, Door, GreenRoom and RedRoom.
After the conversion, we applied Depends-Kotlin to analyze entities and entity relations of ‘maze-kotlin’ and ‘maze-kotlin-java’. We compared the results with those of ‘maze-java’, which is shown in Table 2. The upper part of this table shows the extracted entities at various granularities: Package, File, Class/Interface, Method Member, Java Field and Kotlin Property. The lower part displays the extracted dependency relations at the file level.
Table 2. Entity and relation extracted in maze game.
Entity maze-java maze-kotlin maze-kotlin-java
Package 2 2 2
File 21 21 21
Class, Interface 23 23 23
Method Member 73 17 69
Java Field 23 - 17
Kotlin Property - 31 6
Dependency Relation
Import 15 8 15
Extend 11 11 11
Call 176 121 163
Use 129 155 124
Parameter 48 41 48
Return 15 14 15
JDependsOnK - - 33
KDependsOnJ - - 9
For entities, the number of Package, File and Class is consistent across ‘maze-java’, ‘maze-kotlin’ and ‘maze-kotlin-java’. This alignment is expected as the general structure remains unchanged during the conversion. The number of Method Member decreased by 4, from 73 in ‘maze-java’ to 69 in ‘maze-kotlin-java’. It further decreased to 17 in ‘maze-kotlin’. The decrease of Method Member is because the getter/setter methods are converted into Kotlin property. This table also shows when 6 Java Field in ‘maze-java’ is converted, we have 6 Kotlin Property in ‘maze-kotlin-java’. In general, the number of Java fields should correspond to the number of converted Kotlin properties. However, the conversion algorithm prefers a property over a function when the method of Java is a getter and cheap to calculate [^1] https://kotlinlang.org/docs/coding-conventions.html#functions-vs-properties. In ‘maze-java’, an abstract class MapSite and its subclasses have a method named getColor, which is converted to as a read-only Kotlin property. Therefore, after the conversion, 8 additional Kotlin properties are generated in ‘maze-kotlin’.
For dependency relations, we observed that the number of extend relation has remained consistent among these three projects. For fully conversion from ‘maze-java’ to ‘maze-kotlin’, the number of import relation decreased from 15 to 8. This reduction is attributed to the fact that the Kotlin compiler can automatically conduct type inference, so there is no need to import the specified type. Furthermore, the number of call relations decreased by 55, from 176 to 121. This decrease can be explained by the conversion process, where Java methods are treated as Kotlin properties in ‘maze-kotlin’. Consequently, in ‘maze-kotlin’, accessing these Kotlin properties is categorized to the use relation. This also accounts for the increase in the number of use relations in ‘maze-kotlin’. The decrease in the number of parameter relation is also related to Kotlin directly replacing Java’s getter and setter methods with properties. This means that explicit method calls for accessing parameters are reduced in Kotlin. The number of return relation has remained consistent during the conversion process. This simplifies the process and allows a smooth transition to the converted code.
Our tool also detected the interactions between Java and Kotlin code: 33 JDependsOnK and 9 KDependsOnJ in ‘maze-kotlin-java’ as shown in Table 2. We manually inspected the code and found that Kotlin class extends Java class 3 times: Kotlin class Door extends Java class Mapsite, Kotlin class GreenRoom and RedRoom extend Java class Room. Besides, Kotlin class Door takes Java class Room as parameter 3 times, directly use it twice, and return it once. This finding confirms the accuracy of Depends-Kotlin in capturing Java-Kotlin interactions.
In summary, Depend-Kotlin demonstrates promise in the accuracy of entity and relation resolution in Kotlin and Kotlin-Java. The results indicate that the converted Kotlin entities can be accurately captured, and most of dependency relations in Kotlin code align with the relations in Java code. Furthermore, our findings suggest that, due to the assumption of the IDEA automatic migration process, dependency relations in Kotlin or Kotlin-Java code can differ from Java code.
3.2. Performance Evaluation
In order to evaluate Depends-Kotlin’s performance, we selected three Kotlin projects of different sizes (RubberPicker [^2] https://github.com/Chrisvin/RubberPicker (Kotlin code only), sqlex [^3] https://github.com/sqlex/sqlex (Kotlin and Java code), and ktor [^4] https://github.com/ktorio/ktor (Kotlin code only)) for analysis. We ran Depends-Kotlin on these three projects with a computer (AMD Ryzen 7 4800H@2.9GHz,8GB RAM), and the results of the analysis by Depends-Kotlin are shown in Table  3.
Table 3. Entity and relations extracted in 3 projects.
Entity RubberPicker sqlex ktor
Package 2 45 255
File 11 231 1922
Class, Interface 16 281 3555
Method Member 139 897 13740
Java Field - 162 -
Kotlin Property 71 133 3040
Dependency Relation
Import 7 300 60
Extend 0 45 476
Call 24 711 2786
Use 74 698 4237
Parameter 8 162 1103
Return 2 131 464
Extension 0 45 390
Delegate 0 0 18
JDependsOnK - 1 -
KDependsOnJ - 343 -
#Total_Relation 115 2436 9534
Consumed Time 8s 39s 581s
From this table, we can observe that in a small Kotlin project like RubberPicker, there are no delegate or extension dependency relations. In contrast, delegate or extension relationships can be identified in a larger Kotlin project such as ktor. In the sqlex project, which includes both Java and Kotlin code, we can find the KDependsOnJ and JDependsOnK dependency relations, indicating Kotlin and Java’s interactions.
Table 3 also shows that, for the RubberPicker project with 16 classes, 139 methods and 115 dependency relations, Depends-Kotlin can resolve it in 8s. For the ktor project with 3,555 classes and 9,534 dependency relations, Depends-Kotlin requires 581 seconds for resolution. For a medium project sqlex with 281 classes and 2,436 dependency relations, the consumed time is 39 seconds. We further calculated the consumed time of each module and identified that the most time-consuming step is due to the Antlr v4 module, which is responsible for parsing files into concrete entities. In future work, we plan to investigate this bottleneck to enhance overall performance.
4. Conclusions
This tool paper proposes a prototype, Depends-Kotlin, which extracts entities and dependency relations in Kotlin code and between Kotlin and Java code. A comparison with a confirmed benchmark demonstrates Depends-Kotlin’s accuracy in capturing dependency relations and invocations in Kotlin(-Java) code. With cross-language development gaining popularity (Li et al., 2021; Mathew and Stolee, 2021; El Arnaoty and Servant, 2024), the analysis of links among different languages is essential for such software systems. To our knowledge, our prototype is the first to detect Kotlin-Java dependency relations based on static code syntax analysis. Our prototype can assist developers and researchers in handling migration process from Java to Kotlin and vice versa. Furthermore, the plug-in architecture of our prototype allows the tool to be extendable to other cross-languages, such as Kotlin-JavaScript, Kotlin-Java-JavaScript, etc. This can also aid in understanding the software systems with the interactions between different programming languages.
The results of the consumed time from three Kotlin(-Java) projects show that Depends-Kotlin can achieve good performance on small or medium-sized projects but not very well on large projects. Our future work plans to focus on improving its performance on large-scale projects.
Appendix A Walkthrough
A.1. Download and Build
Depends-Kotlin can be built and run by following 3 steps.
Step 1.:
Clone repository
Our repository is hosted on GitHub, you can clone the source code with the following command:
git clone https://github.com/XYZboom/depends-kotlin.git
Step 2.:
Compile project
Before compiling the project, make sure you have the Java 17 runtime ready. After that, run the following command at the root of your project to compile it:
on Windows
gradlew.bat build
on Linux
gradlew build
Meanwhile, the -x test option can be used to skip testing. At this point, Gradle Wapper will pull the necessary dependencies and compile them. If everything goes smoothly, you will see an output like this.
Figure 2. Screenshot of Depends-Kotlin’s compilation.
Step 3.:
Generate Runnable Jar
After successful compilation, you can use Gradle’s shadow plugin to package it as an executable jar using the following command:
on Windows
gradlew.bat shadowJar
on Linux
gradlew shadowJar
You will see the following output and find an executable jar package with the suffix *. all.jar in the projectRoot/build/libs folder.
Figure 3. Screenshot of Depends-Kotlin’s packaging.
A.2. Run from Command Line
Depends-Kotlin is based on the JVM platform and requires at least Java 17. The input of Depends-Kotlin is the source folder path of a particular project. The full command line arguments are as follows:
Usage: depends <lang> <src> <output> [OPTIONS]
Options:
--auto-include Auto include all paths under
the source path
-h, ---help Show this message and exit
-i, ---include The files of searching path
-d, ---dir The output directory
-f, ---format The output format
-g, ---granularity Granularity of dependency
-s, ---strip-leading-path Strip the leading path
-m, ---n-map-files Output DV8 dependency map file
-p, ---namepattern The name path separators
Arguments:
 lang                 The language of project files
 src                  The directory to be analyzed
 output               The output file name
Figure 4. Running screenshot of sqlex.
For example, if we need to analyze a Kotlin(-Java) project “sqlex”, we can execute the command as follows:
java -jar depends-kotlin.jar kotlin ./sqlex result -d ./out
Depends-Kotlin will analyze the source code in the ./sqlex folder and display the running process in the command console as marked in Figure 4. When the analysis is complete, the console will show the consumed time and the analysis results will be output to the specified folder. In this example, the output (analysis results) of the sqlex project can be found in the ./out folder.
Figure 5 displays the results of dependency relations at the file level in the output folder. The variables shows an ordered list of file names. The cells shows a random list of dependency relations and each cell displays the dependency relations between two source files. For example, the red rectangle shows entities in File 141 call and import entities in File 206.
Figure 5. Output screenshot of sqlex’s dependency results.