### v3

# Depends-Kotlin: A Cross-Language Kotlin Dependency Extractor

Abstract.
Since Google introduced Kotlin as an official programming language for developing Android apps in 2017, Kotlin has gained widespread adoption in Android development. However, compared to Java, there is limited support for Kotlin code dependency analysis, which is the foundation to software analysis. To bridge this gap, we develop Depends-Kotlin to extract entities and their dependencies in Kotlin source code. Not only does Depends-Kotlin support extracting entities’ dependencies in Kotlin code, but it can also extract dependency relations between Kotlin and Java. Using three open-source Kotlin-Java mixing projects as our subjects, Depends-Kotlin demonstrates high accuracy and performance in resolving Kotlin-Kotlin and Kotlin-Java dependencies relations. The source code of Depends-Kotlin and the dataset used have been made available at https://github.com/XYZboom/depends-kotlin. We also provide a screencast presenting Depends-Kotlin at https://youtu.be/ZPq8SRhgXzM.
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/xxxxxxx.xxxxxxxconference: ASE’24: the 39th IEEE/ACM International Conference on Automated Software Engineering (ASE ’24); October 27 – November 01, 2024; Sacramento, California, USAbooktitle: Proceedings of the 39th IEEE/ACM International Conference on Automated Software Engineering (ASE ’24), October 27 – November 01, 2024, Sacramento, California, USAisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
The dependency relations among entities in the source code form the foundations of software architecture analysis (Mo et al., 2019; Cui et al., 2019; Xiao et al., 2022). Since Google introduced Kotlin as an official programming language for developing Android apps in 2017, Kotlin has gained widespread adoption in Android development. According to recent empirical studies, a large number of Android apps have been continuously migrated from Java to Kotlin (Ardito et al., 2020; Mateus and Martinez, 2020). However, compared to dependency analysis tools that support Java, such tool support for Kotlin is limited. To our knowledge, many well-known dependency analysis tools that support the Java language, such as Understand (und, [n. d.]), and DV8 (dv8, [n. d.]), currently do not offer support for the Kotlin language.
Kotlin dependency resolution faces two challenges. First, Kotlin is known as “concise, expressive, and designed to be type and null-safe”. It contains a lot of syntactic sugar to ensure these features, thereby increasing the difficulty of resolving entity types and dependencies. Consider the example in Listing 1: Line 1 declares a class named Bar with a property x of type Int (property in Kotlin is similar to field in Java). Line 3 declares a high-level function named calculate that takes a lambda expression as its parameter. As shown in this line, the lambda takes in a Bar type (called receiver type in Kotlin), and returns an Int. Line 5 declares another class named Foo, also with a property x of type Int. Line 6 declares calculateInFoo, which is a member function of the Foo class. It invokes the calculate function in Line 7. Since the lambda parameter takes type Bar, the add function invoked in the lambda accesses the x property of the current Bar instance, not the Foo instance. The use of lambdas with receiver types as input allows concise syntax but it increases the difficulty of dependency resolution. In order to resolve this dependency (calculateInFoo use Bar.x), we need to locate calculate’s parameters and further trace them down to the correct type and properties. Furthermore, Kotlin has its own dependency types, such as delegate and extension, which lacks support from existing tools..
1class Bar(val x: Int)
2// Declaration of class Bar with the property x
3fun calculate(param: Bar.() -> Int) {}
4/* Function ’calculate’ takes a lambda with a receiver type Bar as its parameter.*/
5class Foo(val x: Int) {
6 fun calculateInFoo() {
7 calculate { add(x) }
8 // x in add(x) here is actually Bar.x
9 }
10}
Listing 1: Kotlin syntax sugar example
Second, Kotlin is designed to be fully interoperable with Java and can run on the JVM, which makes it easy to continuously migrate Java code to Kotlin. Listing 2 includes a Kotlin class BarKotlin and a Java class FooJava. Lines 3-5 demonstrate that the Java class FooJava accepts and interacts with the Kotlin class BarKotlin by invoking a getX() method of the Kotlin class. However, in the Kotlin code the getter method is implicitly generated by JVM for the class property. If we analyze only static code, such dependencies cannot be recognized. Since a significant percentage of Android apps involve both Java and Kotlin code (Ardito et al., 2020; Mateus and Martinez, 2020), a tool for resolving Kotlin dependencies should not only address dependency relations in Kotlin but also handle dependencies between Java and Kotlin.
1class BarKotlin(val x: Int)
2public class FooJava {
3 public static void func(BarKotlin bar) {
4 System.out.println(bar.getX());
5 }
6}
Listing 2: Java-Kotlin implicit invocation
To tackle these two challenges, we propose Depends-Kotlin, a cross-language Kotlin dependency extractor. Depends-Kotlin is based on the Depends framework, an open-source project designed for code dependency analysis (Jin et al., 2019). We enhance Depends to support the Kotlin language by performing multi-round inference for getting Kotlin’s specific types and dependency relations. Additionally, we address cross-dependency relations between Kotlin and Java by refactoring the architecture of the original Depends framework and filling in the implicit code of Kotlin properties, which is handled by JVM and not shown in the source code. Depends-Kotlin is evaluated on three open-source projects containing both Kotlin and Java code and demonstrates its ability to accurately and efficiently extract Kotlin-Java and Kotlin-Kotlin dependencies.
Figure 1. Plug-in architecture of Depends-Kotlin
2. Methodology
2.1. Framework
Figure 1 shows the architecture of Depends-Kotlin. We adopted a plug-in architecture (Wolfinger et al., 2006) and decomposed the original Depends architecture into two parts: Depends-Core (the main framework) and Depends-Lang (plug-ins). Depends-Lang manages the entity parsing for specific languages, while Depends-Core handles the resolution of dependencies for the parsed entities. This architecture refactoring decision aims to enable the handling of cross-language dependencies, which is not supported by the original Depends architecture. To this end, we added in Depends-Core a Language Register module, which leverages Java Service Provider Interface (SPI) mechanism and allows registering different programming languages simultaneously. For a specific language, its language processor needs to extend the interface and has been registered in the SPI file. We also modified the existing Relation Resolver module in Depends-Core to generate the implicit code, specifically the getter/setter code for Kotlin properties mentioned in Section  1. This modification helps the detection of Java-Kotlin interactions.
In order to process Kotlin’s unique syntax and resolve specific dependency relations, we modified the existing Expression Analyser and Relation Resolver modules in Depends-Core. Specifically, We analyzed Kotlin’s extension functions and properties in Expression Analyser. We also identified the scope of Kotlin’s functions by adding related code as the context in Expression Analyser. Furthermore, we adapted Relation Resolver to handle dependencies with built-in types and new unique dependency relations in Kotlin. For example, as method members of built-in types in Java always return built-in types, the original Depends chose to focus on the dependencies of analyzed source files and ignored built-in types. However, due to Kotlin extension functions, built-in types can return any types. This can greatly impact dependencies in the analyzed source code. Consequently we modified the Relation Resolver module to resolve such extension dependency relation.
For Kotlin’s entity parsing, we adopted the same logic of Java-parser in the original Depends framework. We used Antlr v4 to generate an Abstract Syntax Tree (AST) parser from a Kotlin grammar, which is provided on the official Kotlin website. This module is named as Kotlin Processor as shown in Figure 1. Subsequently, the Kotlin EntityTree Builder module store parsed AST nodes into concrete entities, such as files, packages, types, expressions, functions, and properties. In this step, some types in expressions can be resolved in advance, which can reduce subsequent workload. As previously mentioned, the parser module of a specific language can work as a plug-in to the Depends-Core framework, enabling the analysis of other languages in the future.
2.2. Entity and Dependency Relation Resolution
Table 1. Dependency Relations supported by Depends-Kotlin
Relation Description
Import a file imports another class, enum, static method
Contain a class holds another class’s object as field
Extend a class extends a parent class
Implement a class implements an interface
Call an expression invokes another method
Create an expression in a method create an object
Cast an expression does a cast to a type
Annotation an entity uses an annotation
Use a method access a variable in its scope
Parameter a method use another type as its parameter
Return a method returns another type
Delegate a class delegates another class
Extension a method is an extension of a class
Note: The first 11 dependency relations can be observed between Java and Kotlin entities. The last two are exclusive to Kotlin-Kotlin and Kotlin-Java.
Table 2. Dependency Relations Extracted by Depends-Kotlin
DSU-Sideloader Flap RootEncoder Total
K-J J-K K-K J-J K-J J-K K-K J-J K-J J-K K-K J-J
Import 19 0 170 1 3 0 170 1 102 190 913 193 1762
Contain 6 0 34 5 4 0 46 63 10 256 318 73 815
Extend 1 0 4 0 0 0 20 2 6 1 101 75 210
Implement 1 0 0 7 0 0 36 1 9 4 8 3 69
Call 50 0 450 81 3 0 370 479 62 423 1278 1316 4512
Create 1 0 31 2 3 0 94 25 59 71 489 57 832
Cast 0 0 0 1 5 0 10 15 0 8 48 9 96
Annotation 0 0 0 0 0 0 0 2 0 0 0 0 2
Use 70 0 703 187 9 0 713 1093 141 203 3331 4317 10767
Parameter 1 0 48 9 0 0 40 46 29 121 238 145 677
Return 3 0 36 16 0 0 71 24 2 30 263 30 475
Delegate 0 - 0 - 0 - 2 - 0 - 0 - 2
Extension 0 - 0 - 0 - 6 - 0 - 3 - 9
Total 152 0 1476 309 27 0 1578 1751 420 1307 6990 6218
Figure 2. J-K/K-J Dependency visualization(orange denotes K depends on J; green denotes J depends on K)
We followed the same logic of the original Depends framework and performed multi-round inference for getting entities and dependency relations. An entity is represented by a tuple (id, name, entityType, context) and a dependency relation is represented by (sourceId, targetId, dependencyType, weight). For simple dependency relations, such as extend, function parameter, function return and delegate, such dependency relations between two entities can be collected directly when analyzing expressions. For dependency relations, such as member access and member function call on expressions of unknown types, we conducted type inference. The basic idea is to deduct types from what is known to what is unknown. Any operation on an expression with a known type, such as a member function call, will yield the next expression with a known type. This process is recursively repeated until the entire expression’s type inference is completed.
An additional search is required to resolve Kotlin extension functions during this process, as the scope of the expression containing extension functions needs to be identified. While traversing the syntax tree, if a function has a receiver type, it is marked as an extension function and the extended type is the receiver type. After the type system processing is complete, extension functions can be located based on their marks during the traversal of the entity tree. The extended types and the extension relationship of the function can then be recorded.
Our tool framework also automatically generates implicit code for Kotlin, and it distinguishes Java and Kotlin entities by assigning them different labels. When resolving expressions involving Kotlin-Java interactions, such cross-language relations can be easily detected. Table 1 lists the dependency relations currently supported by Depends-Kotlin. Take Listing 2 for example, a Java expression is calling a Kotlin method getX() so a Call relation is extracted with Java source and Kotlin target information.
3. Evaluation
Subjects: Our subjects are DSU-Sideloader (1259 Stars, 94.1% Kotlin, 4.5% Java, 7k LOC), Flap (286 Stars, 52.4% Kotlin, 47.6% Java, 16k LOC) and RootEncoder (738 Stars, 55.8% Kotlin, 42.1% Java, 51k LOC). We chose these projects because they have a high number of stars and range in size from 7k to 51k LOC, allowing us to test our tool’s performance. Additionally, the Kotlin ratio varies among these three projects, from 94.1% in DSU-Sideloader to 55.8% in RootEncoder, which helps us test Depends-Kotlin’s ability to handle Kotlin-Kotlin and Kotlin-Java dependencies.
Run Depends-Kotlin: Following Figure 1’s architecture, Depends-Kotlin was implemented in Kotlin (the new Kotlin-Parser module) and Java (the legacy Depends-Core and Java-Parser modules). The README in the provided GitHub link specifies how to build and run this tool. Essentially, it takes the source file folder as input and outputs a JSON file with entities and dependency relations.
Results: Table 2 presents the extracted dependencies by Depends-Kotlin. The 1st column shows the dependency relations supported by our tool. The 2nd columns list the number of extracted dependencies in DSU-Sideloader, with each sub-column representing dependency relations from a particular source (Java or Kotlin) to a particular destination (Java or Kotlin). For example, 19 in the K-J column under DSU-Sideloader and the Import row means there are 19 instances where Kotlin files import a Java class, enum, or static method. 423 in the J-K column under RootEncoder and the Call row means there are 423 instances where an expression in Java code invokes a Kotlin method in RootEncoder.
Table 2 shows that all 13 dependency relations can be extracted from these three projects, with Use, Call, and Import being the top three most frequent. We also observed that Delegate and Extension have only 2 and 9 instances, respectively, indicating that these two new syntax features are not widely used in these three projects. The last row shows the total number of four kinds of relations in each project. There are 1476 K-K and 309 J-J relations in DSU-Sideloader with 94.1% Kotlin code, compared to 6990 K-K and 6218 J-J relations in RootEncoder with 55.8% Kotlin code. It makes sense that a higher ratio of Kotlin code tends to result in a higher ratio of K-K relations in a project. We also observed a high ratio of K-J (420) and J-K (1307) interactions in RootEncoder and significant K-J relation instances in DSU-Sideloader (152) and Flap (27), which demonstrates our tool’s ability to capture cross-language dependencies. Figure 2 presents the J-K (green arrow) and K-J (orange arrow) dependencies in the three subjects, with each node denoting a source file. As we can see, K-J and J-K dependencies are not limited to specific interfaces, demonstrating good interoperability between Java and Kotlin.
3.1. Accuracy Verification
To our knowledge, there is no available tool to directly extract Kotlin(-Java) dependency relations from source code. Although CodeQL (https://codeql.github.com) and DependencyPlugin (https://github.com/autonomousapps/dependency-analysis-gradle-plugin) can extract dependencies from Kotlin projects, they require compiling the entire project or only provide package-level dependencies, making them unsuitable for comparison with our tool.
Compiler Reference Check: We leveraged JetBrains’ Program Structure Interface (PSI) to conduct a first-round accuracy check. PSI is part of the Java/Kotlin compiler and can find references between elements in programs. If an extracted dependency instance by our tool can be found in references between two elements in PSI, we labeled it as “Found”; otherwise, we labeled it as “NotFound”. We calculate $Precision=\frac{Found\ Instances}{Found\ Instances\ +\ NotFound\ Instances}$ and $Recall=\frac{Found\ Instances}{Dependency\ Instances\ in\ PSI\ References}$ This is a coarse comparison, as PSI does not provide detailed dependency relations and we can only check the existence of extracted dependencies in PSI. Table 3 presents the precision value in our three subjects. Our tool shows promise in resolving dependencies both within the same language and across languages. Specifically, the average precision of J-J and K-K dependencies are 98.2% and 97.9%, respectively, while the average precision of K-J and J-K dependencies are 90.0% and 99.5%, respectively. Generally, the average precision of 97.9% indicates that the capture dependencies by Depends-Kotlin is precise. Due to space limit, the recall table is provided in the dataset link (see the abstract). The average recall across the three subjects is 85.0%, indicating that 85.0% of the dependencies in the ground truth can be correctly captured by Depends-Kotlin.
Manual Check: For each dependency type, we also randomly selected 5 instances from J-J, K-K, J-K, and K-J (if available) in our three subjects, and two authors with four years of Java/Kotlin experience conducted a thorough examination of the instances (the dataset is shared in the GitHub link provided in the abstract). They built the projects in IntelliJ IDEA and independently traced entities to check dependency instances. After completing this, they discussed the results until an agreement was reached. The manual inspection results served as ground truth and were compared with the dependency instances generated by Depends-Kotlin. The comparison shows that our tool can correctly capture 206 out of 213 dependencies (96.7% precision).
In general, both the compiler reference check and the manual check show that Depends-Kotlin can achieve good accuracy in capturing dependency relations.
Table 3. Precision Value of Three Projects
K-J J-K K-K J-J Average
Import 100.0% 100.0% 100.0% 100.0% 100.0%
Contain 100.0% 99.6% 99.4% 100.0% 99.6%
Extend 100.0% 100.0% 100.0% 100.0% 100.0%
Implement 100.0% 100.0% 100.0% 100.0% 100.0%
Call 89.5% 98.8% 98.1% 96.7% 97.4%
Create 100.0% 100.0% 98.2% 95.2% 98.2%
Cast 100.0% 100.0% 100.0% 100.0% 100.0%
Annotation - - - 100.0% 100.0%
Use 78.2% 99.5% 96.7% 98.7% 97.4%
Parameter 100.0% 100.0% 99.7% 99.5% 99.7%
Return 100.0% 100.0% 100.0% 85.7% 97.9%
Delegate - - 100.0% - 100.0%
Extension - - 88.9% - 88.9%
Average 90.0% 99.5% 97.9% 98.2% 97.9%
3.2. Performance Evaluation
Our experiments were conducted on a computer (AMD Ryzen 7 4800H @ 2.9GHz, 8GB RAM), and the performance of our analysis is shown in Table 4. The whole dependency extraction consists of four stages: Source File Parsing, Entity Extraction, Dependency Relation Extraction, and Result Output. We applied instrumentation in the program and calculated each stage’s running time. Due to our Entity Extraction process and Antlr’s Source File Parsing occurring simultaneously, it is difficult to split Source File Parsing and Entity Extraction. We leveraged the IntelliJ Profiler to assess the time ratio consumed by Source File Parsing, then multiplied this ratio by the total running time of the two stages. As shown in Table 4, for project sizes of 7k LOC and 51k LOC, the analysis time ranges from 32.0 seconds to 114.9 seconds. The main functions of this tool —Entity Extraction and Relation Resolution—consumes a reasonable amount of time. The most time consuming stage is Source File Parsing, which is handled by Antlr, taking 87.8%, 77.8% and 78.8% in DSU-Sideloader, Flap and RootEncoder’s total analysis time.
Table 4. Performance of the Four Stages in Three Subjects
DSU-Sideloader Flap RootEncoder
Source File Parsing 28.1s 28.8s 90.5s
Entity Extraction 0.3s 1.1s 3.9s
Relation Resolution 3.3s 6.5s 19.8s
Result Output 0.3s 0.6s 0.7s
Total 32.0s 37.0s 114.9s
4. Conclusions and Future Work
This paper proposes the Depends-Kotlin tool, which can extract dependency relations in Kotlin-Java projects. The evaluation results from three subjects show that Depends-Kotlin can capture Kotlin-Kotlin and Kotlin-Java dependencies accurately and efficiently. With cross-language development gaining popularity (Li et al., 2021; Mathew and Stolee, 2021; El Arnaoty and Servant, 2024), code changes in one language can easily propagate to and impact other languages. Our tool has the potential to assist developers in handling cross-language scenarios, such as Kotlin-Java code smell detection, code migration, and architecture analysis of such systems.
The current version of Depends-Kotlin has several limitations, such as the inability to resolve function calls on variables of generic types or determine the type of variables in lambda expressions (which results in a 85.0% recall), as well as a slow performance for source code parsing by Anltr. We plan to address these limitations in future work, with directions including modifying the Kotlin grammar rules and adopting other parsers with better performance.